# HiCoLoRA: Addressing Context-Prompt Misalignment via Hierarchical Collaborative LoRA for Zero-Shot DST

### v2

@@ B.4 Experiments Implementation Details
The training and validation sets exclude target domain data, while the test set retains only target domain instances. All experiments were conducted on NVIDIA GeForce RTX 5080 GPUs.
Appendix C Case Study
@@ C.1.1 Success Case 1
Dialog Context. We analyze dialog PMUL4648.json (Fig. 6) from the MultiWOZ dataset where a user is seeking information about a restaurant named “saffron brasserie". The dialog involves multiple turns with complex slot-value interactions, including the restaurant name, food type (indian), price range (expensive), area (center).
@@ C.1.2 Success Case 2
Dialog Context. We examine the dialog “8_00066" (Fig. 7) from the SGD dataset, specifically the Trains domain. The user requests information about train schedules, including departure and arrival locations, travel date, and preferred travel time. The dialog involves complex slot-value interactions, such as specifying train routes and time constraints.
@@ C.2.1 Pattern 1: Ambiguous Slot Boundary Cases
Description. These failures occur when slot boundaries are ambiguous or overlapping, making it difficult for the model to distinguish between different slot values or identify the correct slot-value pairs.
Example. In MultiWOZ dialog PMUL4440.json (Fig. 8) involving both restaurant and hotel booking, HiCoLoRA exhibits significant prediction errors. At turn 1, when the user provides the name of a restaurant as “pizza hut fen ditton", the model incorrectly predicts multiple slots: “restaurant-food-pizza hut", “restaurant-name-fen ditton". Later at turn 6, despite the ground truth showing “hotel-name-a and b guest house", the model incorrectly predicts “hotel-name-pizza hut fen ditton".
@@
C.2.2 Pattern 2: Cross-Domain Confusion
@@
Example. In MultiWOZ dialog PMUL3514.json (Fig. 9), HiCoLoRA shows confusion in domain-specific slot value prediction. At turns 3-6, despite the ground truth consistently showing “hotel-name-cityroomz", the model incorrectly predicts “hotel-book day-cityroomz" and “hotel-book people-cityroomz", incorrectly associating the hotel name with booking slots. challenges in semantic entanglement even with disentanglement mechanisms.
@@ C.2.3 Pattern 3: Rare Slot Value Cases
Description. These failures occur when the model encounters rare or unseen slot values that were not adequately represented in the training data. Analysis of the MultiWOZ and SGD datasets reveals that such slots are common: in Attraction, slots like “entrance fee" and “phone" appear in <<10% of dialogs; in Hotel, “stars" and “internet" have fill rates <<20%; in Train, “trainID" appears in <<5% of dialogs. In a zero-shot setting, HiCoLoRA must generalize to both unseen domains and these rare slot values without any domain specific training examples, presenting a significant challenge.
Example. In MultiWOZ dialogs, HiCoLoRA struggles with predicting rare slot values for specific domains. For instance, in attraction domain dialogs, when users request detailed information about “entrance fee" or “address", the model often fails to correctly predict these values. Similarly, in hotel domain dialogs, when users inquire about specific details like “stars" or “internet", the model shows poor performance. In SGD dialogs, similar patterns emerge. For train domain dialogs, HiCoLoRA often fails to predict “trainID" or “price" information, particularly when these values are not explicitly mentioned in the user utterance but are expected as part of the system response.
@@
    Contextual Inference: Properly inferring rare slot values from contextual clues when they are not explicitly mentioned. This is particularly challenging for slots like “trainID" or “reference number" which require the model to generate specific identifiers.

### v3

@@ A.4 Experiments Implementation Details
The training and validation sets exclude target domain data, while the test set retains only target domain instances. All experiments were conducted on NVIDIA GeForce RTX 5080 GPU and Python 3.10.
Appendix B Additional Results and Analyses
B.1 Performance on SGD Dataset
SlotDM-DST 2022 43.9/86.3 36.6/61.4 16.5/62.0 46.7/86.9
T5DST 2021 46.8/N 48.8/N 55.5/N 59.2/N 23.3/N 53.0/N
Prompter 2023 48.4/N 51.5/N 65.3/N 59.2/N 21.9/N 50.8/N
DCC 2023 28.8/N 19.4/N 42.3/N
Table 4 presents the zero-shot performance of HiCoLoRA on the SGD Dataset. Compared to baseline methods and previous state-of-the-art approaches, HiCoLoRA achieves significant improvements across multiple domains.
B.2 Comparison with Contemporary LoRA Methods
Method Year Attr. Hotel Rest. Train Taxi AVG.
HydraLoRA 2024 35.1 18.9 26.3 41.5 65.2 37.4
LoRA-GA 2024 33.8 19.2 24.7 42.8 64.1 36.9
RoSA 2024 36.5 19.6 27.9 43.2 66.8 38.8
Spectral Adapter 2025 37.2 20.1 28.5 43.6 67.3 39.3
HiCoLoRA (Ours) 2025 38.9 20.4 31.0 44.9 68.6 40.8
To situate HiCoLoRA within the evolving landscape of PEFT methods, we compare it against four contemporary LoRA variants: HydraLoRA Tian et al. (2024), LoRA-GA Wang et al. (2024b), RoSA Nikdan et al. (2024), and Spectral Adapter Zhang and Pilanci (2025). As shown in Table 5, HiCoLoRA achieves the highest average JGA, outperforming all baselines in nearly all domains. This superiority is not merely incremental; it stems from fundamental architectural and semantic distinctions that address the core challenges of zs-DST.
Knowledge Preservation and Transfer: LoRA-GA improves the alignment of the gradient during initialization to accelerate convergence but does not explicitly modulate the singular values to align with the specific semantics of the task. HiCoLoRA’s SemSVD-Init not only preserves pre-trained knowledge, but also amplifies singular components relevant to domain-slot structures, effectively mitigating catastrophic forgetting and enhancing zero-shot generalization, particularly for rare slots such as hotel-stars.
HiCoLoRA addresses the unique challenges of zs-DST: cross-layer misalignment, semantic conflation, and knowledge distortion. While other LoRA variants offer general-purpose efficiency, HiCoLoRA provides a domain-aware and layer-conscious design that is essential for robust zero-shot transfer in TODs.
B.3 Scalability Analysis: Generalization Across Model Scales
Method Year Base Model Attr. Hotel Rest. Train Taxi AVG.
ChatGPT-zsTOD 2023 ChatGPT (GPT-3.5) 52.7 42.0 60.8 70.9 55.8 56.4
ChatGPT-zsTOD 2023 ChatGPT (GPT-3.5) 67.2 37.6 67.3 74.4 60.1 61.3
D0T 2024 LLAMA2-13B 63.1 43.8 60.8 48.8 64.7 56.2
MoPE 2024 ChatGLM-6B 60.4 34.1 64.0 71.3 55.9 57.1
FnCTOD 2024 ChatGPT (GPT-4) 58.8 45.2 69.5 76.4 63.2 62.6
FnCTOD 2024 LLAMA2-13B 62.2 46.8 60.9 67.5 60.3 59.5
Multi-User 2025 GPT-4o 56.8 46.0 61.9 69.3 55.1 57.8
HiCoLoRA 2025 LLAMA2-13B 62.0 42.0 61.0 65.0 69.0 60.0
HiCoLoRA 2025 Qwen2.5-14B-Instruct 64.0 44.0 63.0 68.0 71.0 62.0
To rigorously assess the scalability and architectural generality of HiCoLoRA, we extend our evaluation to LLM, comparing against contemporary LLM-based zs-DST methods, including ChatGPT-zsTOD Heck et al. (2023), D0T Finch and Choi (2024), MoPE Tang et al. (2024), FnCTOD Li et al. (2024) and Multi-User Song et al. (2025). As shown in Table 6, HiCoLoRA achieves competitive performance when deployed in LLAMA2-13B and Qwen2.5-14B-Instruct, with an average JGA of 62.0% in the latter, only marginally below FnCTOD with GPT-4 (62.6%) and significantly outperforms other baselines based on LLM.
Architectural Generalization Beyond Scale. The consistent performance of HiCoLoRA in both both small (T5-small, 60M) and large (13B-14B) models underscores a key insight: its hierarchical adaptation mechanism is scale-agnostic. The efficacy of HiCoLoRA stems from its structured semantic alignment decomposition, which addresses cross-layer coordination (RQ1), domain-slot disentanglement (RQ2), and knowledge preservation (RQ3) through explicit inductive biases. This allows it to be generalized effectively even when applied to larger models without architecture-specific modifications.
B.4 Extended Comparison with FnCTOD
Since FnCTOD (Li et al., 2024) achieves comparable performance to HiCoLoRA under the same LLaMA2-13B backbone, we conduct a detailed comparison to highlight their differences in experimental setup and efficiency. A thorough examination of FnCTOD’s experimental configuration reveals several deviations from a strict zero-shot setting.
FnCTOD uses a carefully curated dataset of 7,200 dialogs across 36 domains (including SGD, CamRest676, MSR-E2E, TaskMaster, and WOZ), which include domains overlapping with MultiWOZ test domains. This violates the strict zero-shot learning premise. In contrast, HiCoLoRA uses only 4,625-7,684 samples from 4 domains in MultiWOZ, with one domain excluded during training to ensure a strict zero-shot setting. To ensure a fair comparison, we conducted an additional experiment by training FnCTOD on the FnCTOD dataset. The results, summarized in Table 7, demonstrate that FnCTOD achieves superior performance while maintaining significantly higher efficiency.
Method Attr. Hotel Rest. Train Taxi AVG. Relative Change
FnCTOD (Fine-tuned LLaMA2-13B) 62.2 46.8 60.3 60.9 67.5 59.5 -0.8
FnCTOD (No FT LLaMA2-13B) 49.8 29.5 48.9 53.6 64.7 49.3 -21.1
FnCTOD (GPT-4 SOTA) 58.8 45.2 63.2 69.5 76.4 62.6 +4.2
HiCoLoRA (LLaMA2-13B) 62.0 42.0 61.0 65.0 69.0 60.0
HiCoLoRA (FnCTOD Dataset) 62.8 49.2 63.9 70.3 69.4 63.1 +5.2
Table 7: Performance comparison between HiCoLoRA and FnCTOD under different settings on MultiWOZ (JGA %).
Beyond the fundamental discrepancy in training data composition, our comparative analysis reveals several critical distinctions that underscore HiCoLoRA’s methodological rigor and practical efficiency: (1) When trained on identical data, HiCoLoRA achieves a JGA of 63.1, surpassing FnCTOD by 6.1% and even exceeding GPT-4-based FnCTOD by 0.5 JGA points; (2) HiCoLoRA maintains superior inference efficiency, requiring only a single LLM call with 16 token prompts versus FnCTOD’s dual invocations and larger than 1200 token inputs; (3) While FnCTOD(without fine-tune) employs 5 few-shot examples in its zero-shot configuration (achieving only 49.3 JGA), HiCoLoRA operates under strict zero-shot conditions to attain 60.0 JGA; (4) FnCTOD’s incorporation of detailed schema descriptions deviates from minimal prompt principles, whereas HiCoLoRA relies solely on its hierarchical adaptation mechanism; (5) Architecturally, HiCoLoRA achieves competitive performance through semantic aware initialization and efficient parameter updates, avoiding the computational overhead of prompt heavy approaches.
This comparative analysis demonstrates that FnCTOD not only achieves state-of-the-art performance under strict zero-shot settings but also offers superior efficiency and scalability compared to prompt heavy LLM-based approaches. The gains are attributable to its principled hierarchical adaptation, spectral semantic disentanglement, and knowledge preserving initialization mechanisms that are both empirically effective and practically efficient.
B.5 Ablation Study Results
Method Attr. Hotel Rest. Train Taxi AVG. $\Delta$
HiCoLoRA (Full) 38.9 20.4 31.0 44.9 68.6 40.8 -
w/ Swap Hier Strategies 37.2 19.7 22.9 40.2 67.5 37.4 -3.4
w/o Adaptive Fusion 28.9 19.3 20.3 43.0 68.0 35.9 -4.9
w/o Spec Joint Cluster 36.2 19.8 27.5 42.1 63.6 37.8 -3.0
w/ Kaiming Init 34.3 20.4 27.8 40.4 67.5 38.1 -2.7
w/ PiSSA Init 36.5 20.3 29.0 42.5 67.8 38.9 -1.9
w/ MiLoRA Init 34.1 19.9 26.2 38.5 62.9 36.3 -4.5
Table 8: Ablation study on hierarchical architecture, adaptive fusion, spectral clustering, and initialization of HiCoLoRA on MultiWOZ. Attr. and Rest. are abbreviations for Attraction and Restaurant, respectively. The $\Delta$ column shows the absolute performance drop compared to the full model.
Table 8 validates the contributions and necessity of each core component of HiCoLoRA to its overall performance. This validation is conducted by systematically removing or replacing core components, including the hierarchical strategy, adaptive fusion, spectral clustering, and initialization method.
B.6 Comparison with Recent Heterogeneous Methods
To further validate the effectiveness of HiCoLoRA against contemporary approaches with different architectural paradigms, we conducted comparative analyses with two recently proposed state-of-the-art methods: LDST (Feng et al., 2023) and CAPID (Dong et al., 2024).
Comparison with LDST (EMNLP 2023): LDST proposes an Assembled Domain-Slot Instruction Generation approach for DST. This method generates diverse instruction samples by randomly combining different instruction and input templates during fine-tuning, thereby reducing the model’s sensitivity to prompt variations. For example:
We performed comparative experiments on MultiWOZ 2.1 using the LLaMA-7B backbone for both methods. The results demonstrate that HiCoLoRA maintains 1.9% advantage over LDST (57.8 vs. 56.7 Average JGA). This performance gain, coupled with HiCoLoRA’s parameter efficient design, further validates the effectiveness of our hierarchical collaborative architecture in capturing complex dialog state dependencies. Comparison with CAPID (EMNLP 2024): CAPID proposes Context-aware Auto-prompting and Instruction-following Contrastive Decoding. This approach employs a two stage framework where a context-aware slot query generation method via auto-prompting which initially using GPT-4, aligns the gap between source and target domains. The generated prompts are used to train a T5-base student model to independently produce context-aware slot queries. During inference, the fine-tuned T5-base student model first generates the prompt, which is then used by the trained DST model (T5-base or T5-small) to predict slot values.
Method Configuration Attr. Hotel Rest. Train Taxi AVG.
CAPID T5-base + T5-base 40.9 43.5 37.1 49.5 87.1 50.1
CAPID T5-base + T5-small 33.3 31.1 31.6 34.3 65.4 40.7
HiCoLoRA (Ours) T5-small 38.9 20.4 31.0 44.9 68.6 40.8
Table 9: Comparison with CAPID on MultiWOZ 2.1
We compared HiCoLoRA with CAPID under different model configurations on MultiWOZ 2.1 (Table 9). HiCoLoRA shows a marginal advantage of 0.1% in Average JGA over the CAPID configuration (T5-base + T5-small). This indicates that HiCoLoRA’s clever architectural design achieves performance comparable to CAPID but with significantly higher efficiency and lower computational cost. Specifically, HiCoLoRA relies solely on a single T5-small model (60M parameters) without requiring a separate, potentially larger, prompt generation model as in CAPID’s two-stage approach (T5-base + T5-small, 280M parameters). Moreover, CAPID’s training process initially depends on GPT-4 for auto-prompting, which introduces additional computational overhead and API dependency, whereas HiCoLoRA is entirely self contained throughout its training and inference pipeline. Discussion: HiCoLoRA demonstrates distinct advantages over contemporary approaches. It surpasses the architectural efficiency of full fine-tuning methods like LDST through parameter effective LoRA adaptation, streamlines the multi-stage inference pipeline characteristic of CAPID via a unified hierarchical model, and offers enhanced scalability by natively accommodating multi-domain dialogs without external dependencies. This positions HiCoLoRA as an optimally balanced solution, delivering robust performance with markedly greater practical efficiency for dialog state tracking.
B.7 Generalization Analysis
To rigorously evaluate HiCoLoRA’s robustness and generalization capability in challenging scenarios, we conducted comprehensive cross dataset and cross domain experiments that simulate real world distribution shifts and semantic sparsity conditions. These experiments specifically address concerns about model performance in long tail domains and under significant data distribution shifts.
B.7.1 Cross Dataset Evaluation
We performed extensive cross dataset evaluations to test HiCoLoRA’s ability to generalize across different data distributions and domain structures.
MultiWOZ to SGD Transfer: Trained exclusively on all MultiWOZ domains and evaluated on the complete SGD test set, requiring adaptation to SGD’s broader and unfamiliar service domains. As shown in Table 10, under this challenging setup, HiCoLoRA maintained an average JGA of 47.6%, representing only a 5.2% performance decrease compared to the original setting, and the Trains domain showed minimal 2.0% decline. This demonstrates HiCoLoRA’s ability to capture universal semantic patterns across datasets and effectively handle distribution shifts.
Experiment Buses Events Flights Media Messaging Music Payment Trains AVG.
HiCoLoRA (Original) 54.0 55.1 30.7 75.9 67.7 35.8 26.7 55.8 50.2
HiCoLoRA (MultiWOZ→SGD) 52.4 51.8 29.2 70.6 63.6 33.7 24.8 54.7 47.6
Table 10: Cross dataset generalization performance (JGA %) from MultiWOZ to SGD
SGD to MultiWOZ Transfer: Trained on SGD domains and evaluated on MultiWOZ, testing transfer from diverse but shallower domains to more complex dialog structures. As shown in Table 11, when transferring from diverse but shallower SGD domains to the more complex MultiWOZ, HiCoLoRA maintained an average JGA of 38.8%, a decrease of only 4.9% from the original performance. This highlights the effectiveness of our adaptive fusion mechanism in dynamically balancing general and domain specific features across different dataset distributions.
Experiment Attr. Hotel Rest. Train Taxi AVG.
HiCoLoRA (Original) 38.9 20.4 31.0 44.9 68.6 40.8
HiCoLoRA (SGD→MultiWOZ) 37.0 19.0 30.4 43.1 64.5 38.8
Table 11: Cross dataset generalization performance (JGA %) from SGD to MultiWOZ
B.7.2 Low Semantic Overlap Transfer
To validate the model’s performance in data sparse and semantically unique long tail domains, we conducted a specialized Low Semantic Overlap Transfer experiment. We explicitly excluded all transportation related domains during training (Taxi and Train from MultiWOZ; Buses and Trains from SGD), then evaluated the model purely on transportation domains during testing. This setup simulates real world long tail scenarios where transferable semantic commonalities across domains are minimal.
Experiment Taxi (MultiWOZ) Train (MultiWOZ) Buses (SGD) Trains (SGD) AVG.
HiCoLoRA (Original) 68.6 44.9 54.0 55.8 55.8
HiCoLoRA (Cross-Dataset/Domain) 62.8 38.8 49.7 51.3 50.7
Table 12: Low semantic overlap transfer performance (JGA %) in transportation domains
Under this extreme setting with zero transportation domains in training, HiCoLoRA achieved an average JGA of 50.7% in transportation domains, a decrease of 9.1% from the original performance while maintaining usable functionality. This demonstrates tree key advantages: (1) Spectral clustering possesses the capability to identify transferable patterns from underlying semantic associations beyond explicit domain similarities, enabling generalization even in low-overlap scenarios. (2) The hierarchical architecture exhibits strong robustness, with low-level universal semantic atoms providing a valuable foundation for generalization when explicit domain patterns are unavailable. (3) The adaptive fusion mechanism offers dynamic flexibility, adjusting feature weights based on domain characteristics to avoid over reliance on specific domain patterns and maintain performance under distribution shifts.
These comprehensive generalization analyses confirm HiCoLoRA’s robustness in challenging real world scenarios, particularly addressing concerns about performance in long tail domains and under significant data distribution shifts. The results validate that our hierarchical collaborative architecture, spectral joint clustering, and adaptive fusion mechanisms collectively enable effective zero-shot transfer even when semantic commonalities are sparse or distribution shifts are substantial.
B.8 Inference Efficiency Analysis
HiCoLoRA addresses potential latency concerns in multi-branch designs through a precomputation and merging strategy, ensuring inference efficiency comparable to standard LoRA. We first detail the strategy and then present benchmarking results.
B.8.1 Precomputation and Merging Strategy
UniRep-LoRA Simplicity: The UniRep-LoRA module (Eq. 1) maintains a single set of $\boldsymbol{A}_{ur}$ and $\boldsymbol{B}_{ur}$ matrices throughout.
Heuristic Grouping: For lower layers employing heuristic grouping (Eq. 2), only a single optimal pair $\boldsymbol{A}_{sa}^{*}$ and $\boldsymbol{B}_{sa}^{*}$ is selected, requiring just one matrix multiplication per forward pass.
Full Collaboration: During training, SemAdapt-LoRA employs multiple $\boldsymbol{A}_{sa}^{m}$ and $\boldsymbol{B}_{sa}^{n}$ matrices to enable fine-grained semantic adaptation. During inference, we precompute the collective low-rank contribution of all matrix pairs, and the SemAdapt-LoRA output in full collaboration layers (Eq. 3) can be reorganized by computing aggregated matrices:
$\boldsymbol{A}_{\text{total}}=\sum_{m=1}^{M}\boldsymbol{A}_{sa}^{m},\quad\boldsymbol{B}_{\text{total}}=\sum_{n=1}^{N}\boldsymbol{B}_{sa}^{n}$ (12)
yielding the equivalent computation:
$\boldsymbol{h}_{sa}=\boldsymbol{W}_{0}^{h}\boldsymbol{x}_{sa}+\boldsymbol{B}_{\text{total}}\boldsymbol{A}_{\text{total}}\boldsymbol{x}_{sa}$ (13)
This transformation reduces the computational overhead from $O(M\cdot N)$ matrix multiplications to merely two matrix multiplications, identical to standard LoRA.
The matrix additions involved in precomputation $O(r\cdot d)$ are negligible compared to matrix multiplications $O(d^{2})$. In practice, we precompute all low-rank update terms during model export and absorb them into the base model weights, eliminating any additional inference overhead. Consequently, despite its hierarchical architecture, HiCoLoRA maintains inference latency on par with standard LoRA implementations.
B.8.2 Efficiency Benchmarking Results
Table 13: Inference latency (ms/batch) and GPU memory (MB) comparison across methods. HiCoLoRA adds minimal overhead despite its hierarchical design, remaining efficient for deployment.
Method Latency $\downarrow$ GPU Mem. Base Model (PPTOD-small) $21.2\pm 0.3$ $1,245$ Standard LoRA (r=8) $21.9\pm 0.4$ $1,297$ DualLoRA $22.7\pm 0.4$ $1,315$ HiCoLoRA (Ours) $22.9\pm 0.5$ $1,339$
To empirically validate the inference efficiency of HiCoLoRA, we conduct rigorous benchmarking experiments. All measurements are averaged over 1000 batches after 100 warm-up iterations, using batch size 8 and sequence length 256 on an NVIDIA RTX 5080 GPU. As shown in Table 13, HiCoLoRA introduces only a slight latency increase over others, demonstrating that its hierarchical collaboration is efficiently implemented and does not compromise inference speed.
Appendix C Case Studies
@@ C.1.1 Success Case 1
Dialog Context. We analyze dialog PMUL4648 (Fig. 6) from the MultiWOZ dataset where a user is seeking information about a restaurant named “saffron brasserie”. The dialog involves multiple turns with complex slot-value interactions, including the restaurant name, food type (indian), price range (expensive), area (center).
@@ C.1.2 Success Case 2
Dialog Context. We examine the dialog “8_00066” (Fig. 7) from the SGD dataset, specifically the Trains domain. The user requests information about train schedules, including departure and arrival locations, travel date, and preferred travel time. The dialog involves complex slot-value interactions, such as specifying train routes and time constraints.
@@ C.2.1 Pattern 1: Ambiguous Slot Boundary Cases
Description. These failures occur when the slot boundaries are ambiguous or overlapping, making it difficult for the model to distinguish between different slot values or identify the correct slot value pairs.
Example. In MultiWOZ dialog PMUL4440 (Fig. 8) involving both restaurant and hotel booking, HiCoLoRA exhibits significant prediction errors. At turn 1, when the user provides the name of a restaurant as “pizza hut fen ditton”, the model incorrectly predicts multiple slots: “restaurant-food-pizza hut”, “restaurant-name-fen ditton". Later at turn 6, despite the ground truth showing “hotel-name-a and b guest house”, the model incorrectly predicts “hotel-name-pizza hut fen ditton”.
@@
C.2.2 Pattern 2: Cross Domain Confusion
@@
Example. In MultiWOZ dialog PMUL3514 (Fig. 9), HiCoLoRA shows confusion in domain-specific slot value prediction. At turns 3-6, despite the ground truth consistently showing “hotel-name-cityroomz”, the model incorrectly predicts “hotel-book day-cityroomz” and “hotel-book people-cityroomz”, incorrectly associating the hotel name with booking slots. challenges in semantic entanglement even with disentanglement mechanisms.
@@ C.2.3 Pattern 3: Rare Slot Value Cases
Description. These failures occur when the model encounters rare or unseen slot values that were not adequately represented in the training data. Analysis of the MultiWOZ and SGD datasets reveals that such slots are common: in Attraction, slots like “entrance fee” and “phone” appear in <<10% of dialogs; in Hotel, “stars” and “internet” have fill rates <<20%; in Train, “trainID” appears in <<5% of dialogs. In a zero-shot setting, HiCoLoRA must generalize to both unseen domains and these rare slot values without any domain specific training examples, presenting a significant challenge.
Example. In MultiWOZ dialogs, HiCoLoRA struggles with predicting rare slot values for specific domains. For instance, in attraction domain dialogs, when users request detailed information about “entrance fee” or “address”, the model often fails to correctly predict these values. Similarly, in hotel domain dialogs, when users inquire about specific details like “stars” or “internet”, the model shows poor performance. In SGD dialogs, similar patterns emerge. For train domain dialogs, HiCoLoRA often fails to predict “trainID” or “price” information, particularly when these values are not explicitly mentioned in the user utterance but are expected as part of the system response.
@@
    Contextual Inference: Properly inferring rare slot values from contextual clues when they are not explicitly mentioned. This is particularly challenging for slots like “trainID” or “reference number” that require the model to generate specific identifiers.